OpenStudy (anonymous):

If I'm finding Ci for pi/6, so that it's ((pi/6)+((pi/6)-1) / 2): How do I subtract 1 from pi/6 to get it in terms of pi? The answer should be pi/12

OpenStudy (freckles):

you want to do pi/6-1?

OpenStudy (freckles):

find a common denominator

OpenStudy (freckles):

\[\frac{\pi}{6}-1 \\ \frac{\pi}{6}-\frac{6}{6}\]

OpenStudy (freckles):

you can combine the fractions now

OpenStudy (anonymous):

So I'm getting pi - 6 divided by 6?

OpenStudy (freckles):

if you want to write pi/6-1 as one term yes you get (pi-6)/6

OpenStudy (freckles):

I don't know what Ci means

OpenStudy (freckles):

and it looks like you have another pi/6 up there and something is being divided by 2

OpenStudy (anonymous):

What I'm doing is using the midpoint rule to approximate a definite integral, Ci gives midpoints

OpenStudy (anonymous):

I need to solve the entire thing where I will only have pi/12, which is equal to pi/6 + (pi/6 - 1) all over 2

OpenStudy (anonymous):

I just don't know how to get it in terms of 'nice' pi

OpenStudy (anonymous):

So it's \[\int\limits_{0}^{\pi} \sin x^2 dx\] and solving using the midpoint rule with n = 6

OpenStudy (freckles):

 we need the midpoint of each of these intervals